# From Restless to Contextual: A Thresholding Bandit Approach to Improve Finite-horizon Performance

### v2

@@ Abstract
Online restless bandits extend classic contextual bandits by incorporating state transitions and budget constraints, representing each agent as a Markov Decision Process (MDP). This framework is crucial for finite-horizon strategic resource allocation, optimizing limited costly interventions for long-term benefits. However, learning the underlying MDP for each agent poses a major challenge in finite-horizon settings. To facilitate learning, we reformulate the problem as a scalable budgeted thresholding contextual bandit problem. We carefully integrate the state transitions into the reward design and focus on identifying agents with action benefits exceeding a threshold. We establish the optimality of an oracle greedy solution in a simple two-state setting, and propose an algorithm that achieves minimax optimal constant regret in the online multi-state setting with heterogeneous agents and knowledge of outcomes under no intervention. We numerically show that our algorithm outperforms existing online restless bandit methods, offering significant improvements in finite-horizon performance.
Online Restless Bandit, Contextual Threshold Bandit, Regret Minimization, Finite-horizon Optimization

### v4

# From Restless to Contextual: A Thresholding Bandit Reformulation for Finite-horizon Improvement

@@ Abstract
This paper addresses the poor finite-horizon performance of existing online restless bandit (RB) algorithms, which stems from the prohibitive sample complexity of learning a full Markov decision process (MDP) for each agent. We argue that superior finite-horizon performance requires rapid convergence to a high-quality policy. Thus motivated, we introduce a reformulation of online RBs as a budgeted thresholding contextual bandit, which simplifies the learning problem by encoding long-term state transitions into a scalar reward. We prove the first non-asymptotic optimality of an oracle policy for a simplified finite-horizon setting. We propose a practical learning policy under a heterogeneous-agent, multi-state setting, and show that it achieves a sublinear regret, achieving faster convergence than existing methods. This directly translates to higher cumulative reward, as empirically validated by significant gains over state-of-the-art algorithms in large-scale heterogeneous environments. Our work provides a new pathway for achieving practical, sample-efficient learning in finite-horizon RBs.